OpenStudy (anonymous):

The answer is Vb = Va^2. But I want to know why. I'm guessing it's because of Keplers Third Law - Va is the orbital speed of a satellite at a distance x from the centre of the Earth. If the satellite is at a distance 2x from the centre of the Earth and with the speed of Vb, which of the following statements is correct?

OpenStudy (schrodingers_cat):

When I solve this I get Va^2 = Gm/r and Vb^2 = Gm/2r Equating them you get 2(Vb^2) = Va^2 And the answer was Vb = Va^2 ?????

OpenStudy (abb0t):

What answer?